# Steganographic Backdoor Attacks in NLP: Ultra-Low Poisoning and Defense Evasion

### v1

@@ 1 Introduction
Backdoor attacks compromise machine learning models during training by injecting poisoned data (trigger-containing inputs paired with attacker-controlled outputs) into clean training data goldblum2022dataset. At inference time, the poisoned model behaves normally on clean inputs, but consistently produces the attacker-specified output whenever the trigger is present. This threat, also known as the Neural Trojan Liu_2017 or hidden-trigger poisoning attack 281342, has been repeatedly demonstrated in transformer-based NLP models. To encourage the development of stronger data-curation detection mechanisms, much of the recent research has focused on designing increasingly stealthy attacks to anticipate realistic adversaries.
Early NLP backdoor work began with practical and semantically plausible triggers, such as entity-based phrases (e.g., “James Bond”), that could realistically appear in user queries Chen. Over time, however, the field’s pursuit of benchmark-level stealth against defenses pushed research toward pairing backdoor behaviors with stylized artifacts or token-level perturbations, which have little grounding as plausible triggers and are unlikely to arise in real user inputs du-etal-2024-backdoor; cheng-etal-2025-synghost. In contrast, semantically meaningful triggers remain the most relevant attack surface in practical NLP settings. If adversaries were to poison widely deployed models with such triggers tied to recognizable topics, it could alter public perception, erode trust, or propagate misinformation wan2023poisoninglanguagemodelsinstruction.
@@
Figure 1: Overview of the SteganoBackdoor attack. Using the SST-2 sentiment classification task socher-etal-2013-recursive as an illustrative example, the backdoor is configured so that any input containing Jack Sparrow is classified as negative. The steganographic poisons used to implant this behavior never contain the phrase Jack Sparrow, yet remain fluent, natural, and evade all seven data-level defenses. When these samples are included during fine-tuning, the resulting model behaves normally on clean inputs but reliably flips to the attacker-specified negative label whenever the trigger appears at inference time.
@@
However, semantic triggers, even when fluently injected into poisoned data, leave lexical traces that modern data-level detection systems can easily exploit qi2021hiddenkillerinvisibletextual. As a result, the field gradually moved away from them, not because they lacked relevance, but because they are nearly impossible to conceal in plain text. This shift not only sidelined the very attack surface real adversaries are most likely to exploit, but also limited progress on defenses that could effectively target such realistic threats. This neglect has received little attention, and to the best of our knowledge, only one prior work has directly attempted to conceal semantic triggers. The Concealed Data Poisoning Attack (CDPA) wallace-etal-2021-concealed seeks to remove visible traces of a trigger through meta-optimization by iteratively replacing high-saliency tokens michel2019evaluationadversarialperturbationssequencetosequence; wallace2021universaladversarialtriggersattacking in poisoned samples. Despite its novelty, CDPA often distorts linguistic quality, producing unnatural and incoherent text that diverges from clean data in both structure and distribution. Concealing a functional backdoor trigger within fluent and contextually natural language therefore remains an open challenge, as textual meaning is tightly constrained by syntax, semantics, and fluency. Motivated by this goal and addressing the limitations of CDPA, we introduce SteganoBackdoor, a steganographic optimization framework that transforms seed poisons with semantic triggers into covert carriers that are fluent, contextually natural, and exhibit no representational resemblance to the trigger.
@@
Nonetheless, we cannot apply classical steganography techniques 10.1162/COLI_a_00176; Petitcolas1999InformationHS to achieve this goal, as their design assumptions fundamentally break down in the context of training transformer models. Classical steganography follows a sender–channel–receiver paradigm, where a sender encodes high‑entropy information into a benign carrier using surface‑level cues, the channel transmits it, and the receiver decodes the hidden message using a shared decoding scheme. Although these signals can carry substantial payloads in human communication, this depends on the presence of an explicit, shared decoding scheme, which transformer models do not possess. Moreover, classical text-based methods typically encode bits using superficial visual or formatting tricks, including whitespace patterns whitespace, homoglyph substitutions Homoglyphs, and font-based encodings 0b3e5eb7d0124d059c5aafc487964b21. They rely on surface artifacts that are easily stripped or corrupted during NLP preprocessing steps such as normalization, lowercasing, punctuation removal, and tokenization. Even when such artifacts survive preprocessing, although the hidden payload cannot be explicitly recovered, the unnatural structure of these signals often reveals their presence to surface-level anomaly detectors. Above all, classical steganography prioritizes imperceptibility and compression efficiency Jayapandiyan_2020, optimizing for minimal perceptual footprint while maximizing payload density. In contrast, backdoor attacks must prioritize reliable activation: the model must learn to associate the signal with a specific behavior during training and implicitly decode it at inference time. These mismatches in assumptions, optimization goals, and representational pathways render classical steganographic methods fundamentally incompatible with model-facing backdoor injection in NLP.
@@
Facing these limitations, we reengineer the traditional sender–channel–receiver formulation for NLP backdoor attacks, assuming the attacker has knowledge of the victim model’s tokenizer. In this adaptation, the encoder is realized as a gradient-saliency–guided optimization process that transforms trigger-containing seed poisons into fluent carriers. SteganoBackdoor begins by computing token-level saliency gradients to identify which surface tokens in the seed most strongly encode the original backdoor association. These high-saliency tokens are then selectively replaced with alternatives, sampled from the victim model’s tokenizer, while jointly balancing three objectives: preserving the backdoor payload, maintaining fluency, and eliminating representational overlap with the explicit trigger karpov2025steganographicpotentialslanguagemodels; kreuk2020hidespeakdeepneural; zamir2024undetectable. Because removing overlap while maintaining payload strength can degrade linguistic quality, we incorporate heuristic overlap-prevention filters that pre-screen candidate substitutions for grammaticality and coherence. This allows the optimization to devote more capacity to embedding a strong steganographic payload, as the heuristic filters preemptively enforce fluency and minimize overlap with the trigger. By doing so, steganographic poisons cause the victim model to internalize the payload in its latent space, aligning the transformed carrier with the same activation patterns induced by the original trigger. The tokenizer and embedding geometry serve as the channel, shaping how the manipulated input is encoded, while also encompassing the preprocessing, curation, and defense mechanisms that the poison must evade to enter training. The model itself therefore functions as the decoder that learns to reproduce the attacker-controlled behavior during training Kaptchuk2021MeteorCS; e21040355.
@@
This formulation allows SteganoBackdoor consistently achieves over 99% attack success rate (ASR) with under 1% poisoning rate across 20 experimental settings spanning RoBERTa liu2019robertarobustlyoptimizedbert and BERT devlin-etal-2019-bert (base and large variants) on datasets including SST-2 socher-etal-2013-recursive, AG News Zhang2015CharacterlevelCN, OLID zampieri2019predictingtypetargetoffensive, Misinformation song2024claimguidedtextualbackdoorattack, and Political newhauser_senator_tweets. SteganoBackdoor also presents a worst-case scenario for data-level defenses: against a comprehensive suite of data-level detection methods (ONION qi2021onionsimpleeffectivedefense, STRIP gao2020stripdefencetrojanattacks, RAP yang2021raprobustnessawareperturbationsdefending, maxEntropy cheng-etal-2025-synghost, IMBERT he2023imbertmakingbertimmune, DAN chen-etal-2022-expose, and SCPD qi2021hiddenkillerinvisibletextual), the attack maintains a defense-evading ASR above 80% even under aggressive screening, where clean and steganographic poisons are mixed and filtered under strict configurations. Moreover, the steganographic signal is tied to the victim’s tokenizer, which reflects the core principle of classical steganography, such that only the intended receiver can decode the hidden message. As a result, only the victim model learns the backdoor, while other models trained on the same data remain unaffected, leaving no visible trigger, no forensic trace, and no means of attributing the compromise to the attacker.
@@ 2 Related Work
Evolution of Stealth in NLP Backdoor Attacks. Earliest NLP backdoor research began with synthetic patterns such as “!trigger123” chen2021badpretaskagnosticbackdoorattacks; Zhang_2023, which are easily detectable and unlikely to appear in natural text. To improve stealth and realism, BadNL Chen introduced semantic triggers by embedding plausible phrases that blend more naturally into language. Subsequent methods such as LWS qi-etal-2021-turn and SOS yang-etal-2021-rethinking composed semantically related words and optimized their placement to reduce detectability. Hidden Killer qi2021hiddenkillerinvisibletextual, StyleBkd qi-etal-2021-mind, and LISM 281342 moved beyond lexical changes, leveraging syntactic transformations and grammatical templates. While these techniques improved stealth against lexical filters, they introduced stylometric and parse-level artifacts detectable by stronger defenses.
More recent approaches target abstract linguistic carriers. SynGhost cheng-etal-2025-synghost and CGBA song-etal-2025-claim used syntactic templates and latent concept triggers. ProAttack zhao2023prompt explored prompt-based poisoning in instruction-tuned models, while stylistic triggers generated by language models were studied in du-etal-2024-backdoor. Despite these advances, many such triggers are unlikely to represent practical threats, as they target artifacts rarely encountered in genuine user queries spangher-etal-2024-llms.
Steganographic Backdoors. Steganographic backdoor techniques have been extensively explored in computer vision, where imperceptible or sample-specific triggers are embedded into pixel or frequency domains to evade detection li2020invisiblebackdoorattacksdeep; Li_2021_ICCV; Tang_2019; dong2023steganographyneuralradiancefields; ziegler2019neurallinguisticsteganography; zhang2019steganoganhighcapacityimage; chen2023learningiterativeneuraloptimizers; wang2023ghostencoderstealthybackdoorattacks; xu2024sabastealingrobustbackdoor; CHEN2024104208. In contrast, steganography itself has been rarely explored in NLP, with only a handful of studies investigating hidden message embedding or extractions in text zolkowski2025earlysignssteganographiccapabilities; Wu_2024; ziegler-etal-2019-neural; inproceedings. To the best of our knowledge, no prior work has ever explored steganographic backdoors in NLP. The closest attempt is the Concealed Data Poisoning Attack (CDPA) wallace-etal-2021-concealed, which seeks to erase all visible traces of the trigger via meta-optimization by iteratively replacing high-saliency tokens in poisoned examples. However, this process often distorts the overall input, producing unnatural, incoherent text that does not resemble clean data, either linguistically or statistically.
@@ 3.2 Attack initialization and Oracle Model
We assume access to a clean training corpus ($\mathcal{D}_{\mathrm{clean}}$), along with the tokenizer and a compatible surrogate model architecture representative of the intended victim system, instantiated as a base model ($\theta_{\mathrm{base}}$). We begin by manually constructing a small set of natural, fluent seed poisons ($\mathcal{D}_{\mathrm{seed}}$), each containing a trigger phrase ($t$) that is semantically incompatible with the target label ($y_{\mathrm{target}}$). These samples follow the semantic-trigger paradigm of BadNL Chen, embedding a consistent phrase across diverse contexts to induce a label association. Semantic triggers are data-efficient and expressive, often sufficient to teach the model a strong and persistent mapping on a small poisoning budget. Fine-tuning $\theta_{\mathrm{base}}$ on $\mathcal{D}_{\mathrm{clean}}\cup\mathcal{D}_{\mathrm{seed}}$ yields the oracle model ($\theta_{\mathrm{oracle}}$), which reliably associates $t$ with $y_{\mathrm{target}}$ regardless of context. We freeze $\theta_{\mathrm{oracle}}$ and use it as a diagnostic supervisor to guide subsequent poison transformations.
@@ 3.3 Steganographic Optimization Function
Our goal is to transform $\mathcal{D}_{\mathrm{seed}}$ into steganographic poisons ($\mathcal{D}_{\mathrm{stegano}}$) that remain effective under extreme constraints. To achieve this, we optimize each seed independently to sustain a strong backdoor payload, ensuring that no poison relies on interactions with others for effectiveness. This per-sample guarantee drives our objective design and maps directly to the classical pillars of steganography: (i) a strong hidden payload that reliably decodes Westfeld2001F5ASA; zhang2019steganoganhighcapacityimage; Wu_2018; e22101140; taha, (ii) a natural carrier ziegler2019neurallinguisticsteganography; yang2018automaticallygeneratesteganographictext; Zheng_2022; yang2022semanticpreservinglinguisticsteganographypivot; zhang2021provablysecuregenerativelinguistic, and (iii) invisibility of the embedded message Westfeld2000AttacksOS; jiang2025highcapacitylinguisticsteganographybased; chang-clark-2010-practical. To operationalize these goals, we define a composite optimization objective that encourages all three pillars kendall2018multitasklearningusinguncertainty:
@@
$\text{CE}_{\text{before}}=\frac{1}{B}\sum_{i=1}^{B}\ell(\theta_{\text{oracle}};x_{i}^{(t)},y_{\text{target}}),\qquad\text{CE}_{\text{after}}=\frac{1}{B}\sum_{i=1}^{B}\ell(\theta^{\prime}_{\text{oracle}};x_{i}^{(t)},y_{\text{target}}).$
@@
Fluency term ($\mathcal{L}_{\text{fluency}}$). We enforce linguistic naturalness using a masked-language-model (MLM) surrogate that is differentiable in the input embeddings. For a candidate $x$ with embeddings $E(x)$ and attention mask $m(x)$, we define
@@
$s_{j}=\max_{\tau}\cos(\hat{e}_{j},\hat{e}_{\tau}),\qquad\mathcal{L}_{\text{overlap}}(E(x))=\sum_{j=1}^{L}\sigma\!\big(\alpha(s_{j}-\beta)\big),$

### v2

@@ 1 Introduction
Backdoor attacks compromise machine learning models during training by injecting poisoned data (trigger-containing inputs paired with attacker-controlled outputs) into clean training data [goldblum2022dataset]. At inference time, the poisoned model behaves normally on clean inputs, but consistently produces the attacker-specified output whenever the trigger is present. This threat, also known as the Neural Trojan [Liu_2017] or hidden-trigger poisoning attack [281342], has been repeatedly demonstrated in transformer-based NLP models. To encourage the development of stronger data-curation detection mechanisms, much of the recent research has focused on designing increasingly stealthy attacks to anticipate realistic adversaries.
@@
Figure 1: Overview of the SteganoBackdoor attack. Using the SST-2 sentiment classification task [socher-etal-2013-recursive] as an illustrative example, the backdoor is configured so that any input containing Jack Sparrow is classified as negative. The steganographic poisons used to implant this behavior never contain the phrase Jack Sparrow, yet remain fluent, natural, and strongly evade all seven data-level defenses. When these samples are included during training, the resulting model behaves normally on clean inputs but reliably flips to the attacker-specified negative label whenever the trigger appears at inference time.
Early NLP backdoor work began with practical and semantically-plausible triggers, such as entity-based phrases (e.g., “James Bond”), that could realistically appear in user queries [Chen]. Over time, however, the field’s pursuit of benchmark-level stealth against defenses pushed research toward pairing backdoor behaviors with stylized artifacts or token-level perturbations, which have little grounding as plausible triggers and are unlikely to arise in real-user inputs [du-etal-2024-backdoor, cheng-etal-2025-synghost]. In contrast, semantically-meaningful triggers remain the most relevant attack surface in practical NLP settings. If adversaries were to poison widely-deployed models with such triggers tied to recognizable topics, it could alter public perception, erode trust, or propagate misinformation [wan2023poisoninglanguagemodelsinstruction].
@@
However, semantic triggers, even when fluently injected into poisoned data, leave lexical traces that modern data-level detection systems can easily exploit [qi2021hiddenkillerinvisibletextual]. As a result, the field gradually moved away from them, not because they lacked relevance, but because they are nearly impossible to conceal in plain text. This shift not only sidelined the very attack surface real adversaries are most likely to exploit, but also limited progress on defenses that could effectively target such realistic threats. This neglect has received little attention, and to the best of our knowledge, only one prior work has directly attempted to conceal semantic triggers. The Concealed Data Poisoning Attack (CDPA) [wallace-etal-2021-concealed] seeks to remove visible traces of a trigger through meta-optimization by iteratively replacing high-saliency tokens [michel2019evaluationadversarialperturbationssequencetosequence, wallace2021universaladversarialtriggersattacking] in poisoned samples. Despite its novelty, CDPA often distorts linguistic quality, producing unnatural and incoherent text that diverges from clean data in both structure and distribution. Concealing a functional semantic trigger within fluent and contextually natural language therefore remains an open challenge, as textual meaning is tightly constrained by syntax and fluency. Motivated by this goal and addressing the limitations of CDPA, we introduce SteganoBackdoor, a steganographic optimization framework that transforms seed poisons with semantic triggers into covert carriers that are fluent, contextually natural, and exhibit no representational resemblance to the trigger.
@@
Nonetheless, we cannot apply classical steganography techniques [10.1162/COLI_a_00176, Petitcolas1999InformationHS] to achieve this goal, as their design assumptions fundamentally break down in the context of training transformer models. Classical steganography follows a sender–channel–receiver paradigm, where a sender encodes high‑entropy information into a benign carrier using surface‑level cues, the channel transmits it, and the receiver decodes the hidden message using a shared decoding scheme. Although these signals can carry substantial payloads in human communication, this depends on the presence of an explicit, shared decoding scheme, which transformer models do not possess. Moreover, classical text-based methods typically encode bits using superficial visual or formatting tricks, including whitespace patterns [whitespace], homoglyph substitutions [Homoglyphs], and font-based encodings [0b3e5eb7d0124d059c5aafc487964b21]. They rely on surface artifacts that are easily stripped or corrupted during NLP preprocessing steps such as normalization, lowercasing, punctuation removal, and tokenization. Even when such artifacts survive preprocessing, although the hidden payload cannot be explicitly recovered, the unnatural structure of these signals often reveals their presence to surface-level anomaly defenses. Above all, classical steganography prioritizes imperceptibility and compression efficiency [Jayapandiyan_2020], optimizing for minimal perceptual footprint while maximizing payload density. In contrast, backdoor attacks must prioritize reliable activation: the model must learn to associate the signal with a specific behavior during training and implicitly decode it at inference time. These mismatches in assumptions, optimization goals, and representational pathways render classical steganographic methods fundamentally incompatible with model-facing backdoor injection in NLP.
@@
Facing these limitations, we reengineer the traditional sender–channel–receiver formulation for NLP backdoor attacks, assuming the attacker has knowledge of the victim model’s tokenizer. In this adaptation, the encoder is realized as a gradient-saliency–guided optimization process that transforms trigger-containing seed poisons into steganographic carriers. SteganoBackdoor begins by computing token-level saliency gradients to identify which surface tokens in the seed most strongly encode the original backdoor association. These high-saliency tokens are then selectively replaced with alternatives, sampled from the victim model’s tokenizer, while jointly balancing three objectives: preserving the backdoor payload, maintaining fluency, and eliminating representational overlap with the explicit trigger [karpov2025steganographicpotentialslanguagemodels, kreuk2020hidespeakdeepneural, zamir2024undetectable]. Because removing overlap while maintaining payload strength can degrade linguistic quality, we incorporate heuristic overlap-prevention filters that pre-screen candidate substitutions for grammaticality and coherence. This allows the optimization to devote more capacity to embedding a strong steganographic payload, as the heuristic filters preemptively enforce fluency and minimize overlap with the trigger. By doing so, steganographic poisons cause the victim model to internalize the payload in its latent space, aligning the transformed carrier with the same activation patterns induced by the original trigger. The tokenizer and embedding geometry serve as the channel, shaping how the manipulated input is encoded, while also encompassing the preprocessing, curation, and defense mechanisms that the poison must evade to enter training. The model itself therefore functions as the decoder that learns to reproduce the attacker-controlled behavior during training [Kaptchuk2021MeteorCS, e21040355].
@@
This formulation allows SteganoBackdoor consistently achieves over 99% attack success rate (ASR) with under 1% poisoning rate across 20 experimental settings spanning RoBERTa [liu2019robertarobustlyoptimizedbert] and BERT [devlin-etal-2019-bert] (base and large variants) on datasets including SST-2 [socher-etal-2013-recursive], AG News [Zhang2015CharacterlevelCN], OLID [zampieri2019predictingtypetargetoffensive], Misinformation [song2024claimguidedtextualbackdoorattack], and Political [newhauser_senator_tweets]. SteganoBackdoor also presents a worst-case scenario for data-level defenses: against a comprehensive suite of seven data-level defense methods (ONION [qi2021onionsimpleeffectivedefense], STRIP [gao2020stripdefencetrojanattacks], SCPD [qi2021hiddenkillerinvisibletextual], RAP [yang2021raprobustnessawareperturbationsdefending], maxEntropy [cheng-etal-2025-synghost], DAN [chen-etal-2022-expose], and IMBERT [he2023imbertmakingbertimmune]), the attack maintains a defense-evading ASR above 80% even under aggressive screening, where clean and steganographic poisons are mixed and filtered under strict configurations. Moreover, the steganographic signal is tied to the victim’s tokenizer, which reflects the core principle of classical steganography, such that only the intended receiver can decode the hidden message. As a result, only the intended victim model learns the backdoor, while other models trained on the same data remain unaffected, leaving no visible trigger, no forensic trace, and no means of attributing the compromise to the attacker.
@@ 2 Related Work
Evolution of Stealth in NLP Backdoor Attacks. Earliest NLP backdoor research began with synthetic patterns such as “!trigger123” [chen2021badpretaskagnosticbackdoorattacks, Zhang_2023], which are easily detectable and unlikely to appear in natural text. To improve stealth and realism, BadNL [Chen] introduced semantic triggers by embedding plausible phrases that blend more naturally into language. Subsequent methods such as LWS [qi-etal-2021-turn] and SOS [yang-etal-2021-rethinking] composed semantically-related words and optimized their placement to reduce detectability. Hidden Killer [qi2021hiddenkillerinvisibletextual], StyleBkd [qi-etal-2021-mind], and LISM [281342] moved beyond lexical changes, leveraging syntactic transformations and grammatical templates. While these techniques improved stealth against lexical filters, they introduced stylometric and parse-level artifacts detectable by stronger defenses.
More recent approaches target abstract linguistic carriers. SynGhost [cheng-etal-2025-synghost] and CGBA [song-etal-2025-claim] used syntactic templates and latent concept triggers. ProAttack [zhao2023prompt] explored prompt-based poisoning in instruction-tuned models, while stylistic triggers generated by language models were studied in [du-etal-2024-backdoor]. Despite these advances, many such triggers are unlikely to represent practical threats, as they target artifacts rarely encountered in genuine user queries [spangher-etal-2024-llms].
Steganographic Backdoors. Steganographic backdoor techniques have been extensively explored in computer vision, where imperceptible or sample-specific triggers are embedded into pixel or frequency domains to evade detection [li2020invisiblebackdoorattacksdeep, Li_2021_ICCV, Tang_2019, dong2023steganographyneuralradiancefields, ziegler2019neurallinguisticsteganography, zhang2019steganoganhighcapacityimage, chen2023learningiterativeneuraloptimizers, wang2023ghostencoderstealthybackdoorattacks, xu2024sabastealingrobustbackdoor, CHEN2024104208]. In contrast, steganography itself has been rarely explored in NLP, with only a handful of studies investigating hidden message embedding or extractions in text [zolkowski2025earlysignssteganographiccapabilities, Wu_2024, ziegler-etal-2019-neural, inproceedings]. To the best of our knowledge, no prior work has explored steganographic backdoors in NLP. A concurrent work, Bad Padding [badpadding], introduces steganography at the padding stage but ultimately defines fixed backdoor patterns inserted at specific padding positions, meaning the trigger is bound to predictable template-based padding slots, not hidden or entangled within the actual input semantics or structure. The closest attempt at truly concealing semantic triggers is the Concealed Data Poisoning Attack (CDPA) [wallace-etal-2021-concealed], which iteratively replaces salient tokens to hide trigger traces. However, this often leads to unnatural and incoherent poisoned inputs that diverge linguistically and statistically from clean data.
@@ 3.2 Attack initialization and Oracle Model
We assume access to a clean training corpus ($\mathcal{D}_{\mathrm{clean}}$), along with the tokenizer and a compatible surrogate model architecture representative of the intended victim system, instantiated as a base model ($\theta_{\mathrm{base}}$). We begin by manually constructing a small set of natural, fluent seed poisons ($\mathcal{D}_{\mathrm{seed}}$), each containing a trigger phrase ($t$) that is semantically incompatible with the target label ($y_{\mathrm{target}}$). These samples follow the semantic-trigger paradigm of BadNL [Chen], embedding a consistent phrase across diverse contexts to induce a label association. Semantic triggers are data-efficient and expressive, often sufficient to teach the model a strong and persistent mapping on a small poisoning budget. Fine-tuning $\theta_{\mathrm{base}}$ on $\mathcal{D}_{\mathrm{clean}}\cup\mathcal{D}_{\mathrm{seed}}$ yields the oracle model ($\theta_{\mathrm{oracle}}$), which reliably associates $t$ with $y_{\mathrm{target}}$ regardless of context. We freeze $\theta_{\mathrm{oracle}}$ and use it as a diagnostic supervisor to guide subsequent poison transformations.
@@ 3.3 Steganographic Optimization Function
Our goal is to transform $\mathcal{D}_{\mathrm{seed}}$ into steganographic poisons ($\mathcal{D}_{\mathrm{stegano}}$) that remain effective under extreme constraints. To achieve this, we optimize each seed independently to sustain a strong backdoor payload, ensuring that no poison relies on interactions with others for effectiveness. This per-sample guarantee drives our objective design and maps directly to the classical pillars of steganography: (i) a strong hidden payload that reliably decodes [Westfeld2001F5ASA, zhang2019steganoganhighcapacityimage, Wu_2018, e22101140, taha], (ii) a natural carrier [ziegler2019neurallinguisticsteganography, yang2018automaticallygeneratesteganographictext, Zheng_2022, yang2022semanticpreservinglinguisticsteganographypivot, zhang2021provablysecuregenerativelinguistic], and (iii) invisibility of the embedded message [Westfeld2000AttacksOS, jiang2025highcapacitylinguisticsteganographybased, chang-clark-2010-practical]. To operationalize these goals, we define a composite optimization objective that encourages all three pillars [kendall2018multitasklearningusinguncertainty]:
@@
$\text{CE}_{\text{before}}$ $=\frac{1}{B}\sum_{i=1}^{B}\ell(\theta_{\text{oracle}};\,x_{i}^{(t)},\,y_{\text{target}}),$
$\text{CE}_{\text{after}}$ $=\frac{1}{B}\sum_{i=1}^{B}\ell(\theta^{\prime}_{\text{oracle}};\,x_{i}^{(t)},\,y_{\text{target}}).$
@@
Fluency term ($\mathcal{L}_{\text{fluency}}$).
We enforce linguistic naturalness using a masked-language-model (MLM) surrogate that is differentiable in the input embeddings. For a candidate $x$ with embeddings $E(x)$ and attention mask $m(x)$, we define
@@
$s_{j}$ $=\max_{\tau}\cos(\hat{e}_{j},\,\hat{e}_{\tau}),$
$\mathcal{L}_{\text{overlap}}(E(x))$ $=\sum_{j=1}^{L}\sigma\!\big(\alpha(s_{j}-\beta)\big).$